OpenStudy (anonymous):

find the roots of the polynomial square: x^2 + 24 = –11x

OpenStudy (anonymous):

answers a)12,2 b)3,8 c)-8,-3 d)-12,-2

OpenStudy (anonymous):

put all terms to one side of the equation, then factor the equation. 8(3)=24 and 8+3=11 so the factors are 8 and 3 and then the roots would be -8 and -3. So your answer is "c"

OpenStudy (anonymous):

this is a simple quadratic and so you should factor it......faster then the quadratic formula.

OpenStudy (anonymous):

you can use completing the square or the quadratic formula or just factoring :)

OpenStudy (anonymous):

x^2+24=-11x is the same as x^2+11x+24=(x+3)(x+8) and if 0=(x+3)(x+8), well then the roots are x=-3,-8

OpenStudy (anonymous):